Sol Summary – November 24th

Sol 2 Summary Report
November 24, 2017
Julielynn Wong, Crew 183 Mission Commander
Summary Title: A day of continuous 3D printing off-line and off-grid at MDRS.
Mission Status: Nominal.
Sol Activity Summary: We have been measuring the power usage of 3D printing 7 designs (including wind turbine components and a spare part for our drone) on our humanitarian 3D printing system (H3DPS-LM).  Our Crew Engineer is preparing to use our recycler to convert printed plastic items into 3D printer feedstock.
Look Ahead Plan: Preparing for sim on 0700 MST 2017-11-25 (Sol 3) and 0700 MST 2017-11-27 (Sol 5).
Anomalies in work:  Out of sim.
Weather:  Calm, sunny and warm.
Crew Physical Status:  Healthy
EVA: Crew Astronomer, Crew Engineer, HSO, and Greenhab Officer participated in an EVA led by our Crew Geologist.
Reports to be filed:
GreenHab Report
Operations Report
EVA Report
Astronomer Report
Support Requested: None

Science Report – November 24th

Today we took an extended EVA to survey the area. Tomorrow we will perform and EVA in an attempt to acquire 3D scans on specific, marked, specimen in the area as well as attempting to fly our drone over the Martian landscape.

The rest of today was dedicated to designing 3D models for the humanitarian 3D printing experiment. I have completed the designs for the Doctors Without Borders laptop stand and have redesigned the previous Suction Canister project. Both will be printing overnight with results expected immediately afterwards.
The portion of tomorrow thatI I do not use for our EVA will hopefully be used familiarizing myself with the Solar Observation equipment.
I’m excited for our first day of full sim.

Astronomy Report – November 24th

Astronomy Report
Nov. 23rd/2017
Given that the Musk Observatory is still offline (as it probably will be for the remainder of the mission), my Astronomy observation is limited to Solar observations during the day. This is fine, I’m sure I’ll find plenty to keep me occupied with that.
Today we calibrated the 3D printers and have been continuously printing items that we may need to use in remote locations.
Later tonight, much like last night, I plan to perform a non-sim EVA with our Mission Engineer to practice some Astrophotography. Will post pictures in my next report.
Hopefully I get to do some daytime Solar observations tomorrow. I’ll likely spend most of my day figuring out how to operate that lab.
Problems Encountered: The Musk Observatory being offline.

 

Toluwa Fayemi
MDRS Crew 183 Astronomer

EVA Report – November 24th

EVA #2 Report of Sol 2

We started our EVA at 11:50 AM. Jordan, Ashton, Toluwa, and I, each took an ATV, carrying cameras and other supplies in backpacks. We first stopped near the Stream Turnout Rd, and walked about 200 meters to an area with a large amount of dinosaur bones and petrified wood specimens. We spent about an hour in this area before heading north.

We arrived at the Lith Canyon turnout by about 12:20 PM. From here, we walked into the Lith Canyon area, careful not to step in the cactus gardens along the way.

We spent 2 hours here, exploring and photographing this area. There is a feature that is labeled on the map, called “The Pillar”, where the Copernicus Hwy North used to extend to. Although there are many features in this area that could fit this description, we found one particularly prominent column that could be “The Pillar”. We’ll have to investigate this further.

We returned to our ATVs by around 3:00 PM, getting back to the hab by 3:25 PM.

Have a great night Mission Support!

Prepared by Brandon Ferguson

GreenHab Report – November 24th

Sol 2 – 24/11/2017

Green Hab Officer Jordan Bowden

Average temperature: 27.3 C in Greenhouse / 9.6 C Outside

Time 11:00 16:30 19:00
Temperature inside (°C) 46 C 19 C 17 C
Temperature outside (°C) 16 C 7 C 6 C
% Humidity 17% 16% 26 %

Changes to crops:  Everything is growing well. The adult tomato plants are no longer wilted.

Daily water usage for crops: 4 Gallons

Time(s) of watering for crops:  11:00, 16:30

Daily watering and amount of water used:   2 gallons used at each watering, 4 gallons total.

Narrative: Greetings CapCom, things are going well in the Green Hab on Sol 2!

I noticed that it was very hot in the Green Hab when I checked it at 11:00 this morning (46C); I will be sure to open the ventilation tomorrow by 8AM. Otherwise, all crops are healthyl.

I installed lighting timer in the Green Hab today, so the grow lights are now online from 17:00-21:00.

Not much else to report. Outside my duties as Green Hab Officer, I set up our 3D printer server, designed two 3D printable medical devices, and explored Lith Canyon during our EVA.

All the best,

Jordan Bowden

Green Hab Officer

Operations Report – November 24th

Crew 183 Operations Report 24-11-2017

SOL: 2

Name of person filing report: Ashton Stoop

Non-nominal systems:

Power system – SOC

Notes on non-nominal systems:

The power system calculates SOC based on an algorithm, it is not a measured value. This has been causing strange fluctuations on when reading the SOC.

READING FROM 24 NOV 2017

TIME     SOC READINGS

6:32       61%

9:25       69%

10:11     100%

11:40     100%

16:10     88%

18:37     100%

Generator (hours run): 14

Solar— 100%

Diesel –  62%

Propane –  42% volume

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons remaining

Water (trailer) – 100 gallons added to static

Water (static) –   325 gallons

Trailer to Static Pump used –  yes

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  45

Toilet tank emptied: yes

ATV’s Used: Honda, 300, 350.1, 350.2

Oil Added? No

ATV Fuel Used:  5 Gals

# Hours the ATVs were Used today: 3

Notes on ATVs: ATV 350.1 had a bit of trouble starting up upon return from EVA. Once it got going everything seemed to be fine. We’ll keep an eye on this for the next few SOLs.

Deimos rover used:   Yes

Hours: 1

Beginning charge: 100%

Ending charge: 90%

Currently charging: Yes

 

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Spirit rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

 

Opportunity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  NO

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? NO

General notes and comments:

The SOC readings seem off. Looking to further investigate this matter.

Summary of internet:

We hit our internet limit around 20:00.

Summary of suits and radios:

Not used

Summary of Hab operations:

Everything is in good shape. We had some bomb crab biscuits for dinner.

Summary of GreenHab operations:

A new lighting timer has been delivered and installed in the GreenHab.

Summary of ScienceDome operations:

New lights have been delivered. Tomorrow, we need to install lightbulbs and test.

Summary of RAMM operations:

Not used. It is not yet fully operational.

Summary of health and safety issues:

Four carbon monoxide detectors have been ordered.

Questions, concerns and requests to Mission Support:

-Dear mission support, how exactly does the SOC reading on the power system operate? We have had fluctuations in the SOC readings which do not correlate closely with our estimated values of power generation and power consumption (Please refer to SOC Readings table above).

-We have no additional supply of gasoline for the ATVs. All that’s left is what’s in the vehicles. We need fuel ASAP!